### v1

# Enhancing Monotonic Modeling with Spatio-Temporal Adaptive Awareness in Diverse Marketing

Abstract.
In the mobile internet era, the Online Food Ordering Service (OFOS) emerges as an integral component of inclusive finance owing to the convenience it brings to people. OFOS platforms offer dynamic allocation incentives to users and merchants through diverse marketing campaigns to encourage payments while maintaining the platforms’ budget efficiency. Despite significant progress, the marketing domain continues to face two primary challenges: (i) how to allocate a limited budget with greater efficiency, demanding precision in predicting users’ monotonic response (i.e. sensitivity) to incentives, and (ii) ensuring spatio-temporal adaptability and robustness in diverse marketing campaigns across different times and locations. To address these issues, we propose a Constrained Monotonic Adaptive Network (CoMAN) method for spatio-temporal perception within marketing pricing. Specifically, we capture spatio-temporal preferences within attribute features through two foundational spatio-temporal perception modules. To further enhance catching the user sensitivity differentials to incentives across varied times and locations, we design modules for learning spatio-temporal convexity and concavity as well as for expressing sensitivity functions. CoMAN can achieve a more efficient allocation of incentive investments during pricing, thus increasing the conversion rate and orders while maintaining budget efficiency. Extensive offline and online experimental results within our diverse marketing campaigns demonstrate the effectiveness of the proposed approach while outperforming the monotonic state-of-the-art method.
Monotonic Modeling; Online Food Ordering; Marketing Policy
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06submissionid: 744ccs: Information systems Location based services
1. Introduction
With mobile payment services, Online Food Ordering Service (OFOS) platforms such as Ele.me,111www.ele.me.com Meituan 222www.meituan.com and Uber Eats333www.ubereats.com have become critical components of inclusive finance through the convenience offered by real-time provisioning. To encourage mobile payments for food and retail delivery, platforms offer incentives (e.g., coupons, commissions, bonuses) to users through diverse marketing strategies and campaigns. Within these marketing campaigns, the intelligent allocation of incentives to users and merchants represents a critical aspect for platforms in enhancing marketing efficiency.
Users exhibit different response patterns to platform incentives across various marketing campaigns. Overall, the user Conversion Rate (CVR) expresses a monotonic increasing trend to the increment of red packets or discount amounts, while it monotonically decreases about the threshold for coupon. Similarly, CVR exhibits a monotonic decreasing relationship with delivery fees.
Previous work in the marketing domain (Boutilier and Lu, 2016; Beheshti-Kashi et al., 2015) has explored various methods for optimizing incentives under a limited budget. (Ito and Fujimaki, 2017; Zhao et al., 2019; Yu et al., 2021) have approached marketing pricing by employing a paradigm consisting of two components: predicting user response scores and making real-time decisions through linear programming. Importantly, in the takeaway domain, where user preferences are significantly influenced by spatio-temporal factors, relevant studies in the recommendation field (Lin et al., 2022; Du et al., 2023; Lin et al., 2023) have explored the learning and enhanced utilization of spatio-temporal information.
However, previous research in the takeaway marketing domain has overlooked the critical importance of perceiving spatio-temporal information, focusing instead on analyzing users’ sensitivity to incentives in a general context. In reality, users exhibit varying sensitivities to platform incentives when ordering from different merchants at different times and locations. Relying solely on a uniform sensitivity function to assess user response across varying times and locations is flawed. Therefore, it is essential to differentiate descriptions of user sensitivity by spatio-temporal information awareness. Meanwhile, ensuring the model possesses enhanced predictive accuracy and adapts robustly to diverse marketing.
In this paper, we tackle these challenges by proposing Constrained Monotonic Adaptive Network (CoMAN), a novel marketing pricing model utilizing spatiotemporal-enhanced monotonic response. Specifically, our method employs two spatio-temporal aware activation modules to guide the model in learning and understanding spatio-temporal features, thus facilitating a more reasonable capture of spatio-temporal information within the representations. Additionally, our meticulously designed monotonic layer enhances the perception of the spatio-temporal differentiation in users’ monotonic responses to incentives. Concurrently, we introduce the activation function tailored to adaptively learn the incentive sensitivity function of users in the marketing domain. Ultimately, CoMAN achieves a strict continuous monotonic, precise, and smooth fitting of the incentive sensitivity function across different spatio-temporal dimensions. The effectiveness of our proposed methodology is demonstrated through extensive online and offline experiments across diverse marketing campaigns. In summary, our main contributions are as follows:
  • We highlight the primary challenges encountered in the marketing domain and analyze the significance of spatio-temporal preference awareness, while further designing a novel model to address these issues. To the best of our knowledge, our model is the first to systematically model monotonic response by perceiving spatio-temporal information for diverse takeaway marketing.
  • We propose a monotonic layer that utilizes the perception of spatio-temporal information to enhance modeling monotonic responses. This module adaptively learns the concavity, convexity, and parametric expression of user incentive sensitivity functions in diverse marketing campaigns conditioned on different spatio-temporal characteristics.
  • We conceptualize a strictly monotonic Convex Linear Unit (CLU) activation function tailored to diverse marketing. We utilize the activation function to flexibly learn the differences in monotonic function expressions. We demonstrate that this strictly convex function can more effectively approximate user incentive response functions.
2. Related Work
Activation Functions. Activation functions play a pivotal role in neural networks by introducing nonlinear transformations, allowing deep models to capture complex patterns beyond the reach of linear operations(Rumelhart et al., 1986; Neal, 1992; Hochreiter, 1991; Nair and Hinton, 2010). A series of studies like Leaky ReLU (Maas et al., 2013), PReLU (He et al., 2015), and ELU (Clevert et al., 2015) have designed ReLU variants that mitigate this issue by permitting a small, non-zero gradient to negative inputs, thereby mitigating the issue while enhancing the model performance. A series of works by (Shang et al., 2016; Zagoruyko and Komodakis, 2017; Eidnes and Nøkland, 2017) have investigated how to combine the output of the original ReLU with its point reflection within the same neural network layer.
Monotonicity by architecture. Research methods for modeling monotonic response can be categorized into two primary types. The first involves the monotonic structure by construction, such as constrained network weights to be exclusively non-negative or non-positive, guaranteeing the model exhibits either monotonically increasing or decreasing. The other employs regularization to impose monotonicity, networks implement monotonic through the modified loss functions or the heuristic regularization terms during training. Min-Max networks (Sill, 1997; Daniels and Velikova, 2010) utilize monotonic linear embedding and max-min-pooling to endow networks with broad modeling capacity but are challenging to train and practice. Deep Lattice Network (DLN)(You et al., 2017) applies constraints to pairs of lattice parameters but limits the function approximation space and necessitates a substantial number of parameters learned. CMNN (Runje and Shankaranarayana, 2023) can approximate any continuous monotonic function by combining non-saturated and saturated activation functions while lacking the capability to learn adaptively.
Figure 1. Overview of our proposed Constrained Monotonic Adaptive Network (CoMAN) architecture.
Monotonicity by regularization. As mentioned above, enforcing monotonicity during the training can also be achieved by modifying loss function or incorporating regularization terms(Sill and Abu-Mostafa, 1996; Gupta et al., 2019). A. Sivaraman et al.(Sivaraman et al., 2020) introduced a mechanism based on Satisfiability Modulo Theories (SMT) solvers capable of identifying point pairs violating monotonicity constraints (i.e. counterexamples). Liu et al.(Liu et al., 2020) transforms the mathematically verifying monotonicity on arbitrary piece-wise linear (i.e. ReLU) networks into a Mixed Integer Linear Programming (MILP) problem and can be efficiently solved with existing MILP solvers. Both MILP and SMT methodologies, depend on the piecewise-linear property of ReLU to validate the constrained monotonic learning while not being extended to other activation functions such as ELU(Clevert et al., 2015), SELU(Klambauer et al., 2017), GELU(Hendrycks and Gimpel, 2016), etc. Moreover, the computational costs associated with these techniques are significant, frequently necessitating multiple iterations to achieve a rigorously verified monotonic model.
3. Preliminaries
We formulate the marketing problem as a constrained linear programming problem to be solved using the optimization approach. We define the user set as $\mathcal{C}$ and the candidate set of incentive values as $\mathcal{T}$, where $t_{j}\in\mathcal{T}$ represents the $j$-th treatment. Our goal is to find an incentive allocation strategy that optimally allocates incentives to maximize the sum of profits from user response (i.e. CVR) under the constrained budget $\mathcal{B}$. Formally,
(1) $\max_{x_{i,j}}\sum_{i=1}^{|\mathcal{C}|}\sum_{j=1}^{|\mathcal{T}|}x_{i,j}r_{i,j}\$
(2) $s.t.$ $x_{i,j}\in[0,1],for\ i=1,...,|\mathcal{C}|;\ j=1,...,|\mathcal{T}|$
(3) $\sum_{j=1}^{|\mathcal{T}|}x_{i,j}=1,for\ i=1,...,|\mathcal{C}|$
(4) $\frac{\sum_{i=1}^{|\mathcal{C}|}\sum_{j=1}^{|\mathcal{T}|}x_{i,j}t_{j}r_{i,j}}{\sum_{i=1}^{|\mathcal{C}|}\sum_{j=1}^{|\mathcal{T}|}x_{i,j}r_{i,j}}\leq\mathcal{B}$
where $x_{i,j}$ denotes the decision vector indicating whether treatment $t_{j}$ is allocated to user $i$, and $r_{i,j}$ represents the response score by our response model. And $\sum^{|\mathcal{T}|}_{j=1}x_{i,j}t_{j}$ is the incentive value for the user under constraint 3. Furthermore, constraint 4 indicates the limited average incentive to each user. By introducing dual variables $\mu_{i}$ and $\lambda$ corresponding to the aforementioned constraints 3 and 4, we transform the primal problem into its dual counterpart. The convex optimization problem delineated in objective function 1 is addressed through the application of the Lagrangian multiplier method, formally depicted as follows:
(5) $-\sum_{i=1}^{|\mathcal{C}|}\sum_{j=1}^{|\mathcal{T}|}x_{i,j}r_{i,j}+\sum_{i=1}^{|\mathcal{C}|}(\mu_{i}(\sum^{|\mathcal{T}|}_{j=1}x_{i,j}-1))$
$+\lambda(\sum_{i=1}^{|\mathcal{C}|}\sum_{j=1}^{|\mathcal{T}|}x_{i,j}r_{i,j}(t_{j}-\mathcal{B}))$
Rearrange the equation 5, we have:
(6) $\sum_{i=1}^{|\mathcal{C}|}\sum_{j=1}^{|\mathcal{T}|}x_{i,j}[r_{i,j}(\lambda t_{j}-\lambda\mathcal{B}-1)]$
$+\sum_{i=1}^{|\mathcal{C}|}(\mu_{i}(\sum^{|\mathcal{T}|}_{j=1}x_{i,j}-1))$
Given both the objective function and the constraint functions are convex, the optimal $\lambda^{*}$ is obtained by applying the KKT conditions and L-BFGS. From 6, it can be observed that different incentive values are only related to the first term. Consequently, an approximate solution can be derived as follows:
(7) $x_{i,j}=\begin{cases}1,&\text{ if }j=argmin_{j}r_{i,j}(\lambda^{*}t_{j}-\lambda^{*}\mathcal{B}-1)\\ 0,&\text{ otherwise }\end{cases}$
During online pricing decisions, we employ our response model to predict response scores $|\mathcal{T}|$ times based on candidates for each request. Then utilizing the real-time computed value of $\lambda^{*}$ along with decision equation 7 to price the incentive value.
4. Methods
In the marketing campaigns, we design a multi-task model as a pricing model predicting various business objectives such as Conversion Rate (CVR) or CTCVR (Click Through & Conversion Rate) and GMV(Gross Merchandise Volume). Given that the tasks’ predictions lack substantive relevance, there is a risk of encountering the Seesaw Phenomenon during multi-task learning. It may lead one task to improve while another deteriorates. To mitigate this, we utilize the Progressive Layered Extraction (PLE)(Tang et al., 2020) multi-task framework and propose the Constrained Monotonic Adaptive Network (CoMAN).
Overview. The design and modeling of our proposed framework focus on three critical aspects: activating spatio-temporal correlations within the attribute features, modeling monotonic response to incentives in diverse marketing, and enhancing the model’s adaptability to spatio-temporal disparities within the monotonic layer. The overall architecture is depicted in Figure 1.
(a) Four-Parameter Model
(b) Spatio-temporal Attention
Figure 2. Original Four-Parameter Model (FPM) and combined with Spatio-temporal Attention (S-t Attention) fitting.
In our feature extraction process, we segment the period into five intervals: breakfast [04:00–10:00), lunch [10:00–14:00), afternoon tea [14:00–17:00), dinner [17:00–20:00), and midnight snack [20:00–04:00), and we statistic numerical features to each period. We construct temporal features by incorporating information such as the Gregorian and lunar dates, holiday occurrences, and weekdays for each day. Simultaneously, we formulate spatial features from geographical data, encompassing the user’s current location, city ID, district ID, and AOI (Area of Interest) ID. Notably, in certain marketing campaigns different from user-level modeling, we employ regional modeling for AOIs. Therefore, depending on marketing campaigns, we tailor our features to use either users or AOIs (e.g., cross features, historical click sequences).
Spatio-temporal Correlation Activation. Within the Temporal Activation module, spatio-temporal features input to compute the weights of each feature embedding (i.e., User, Shop, Query) $e_{l_{i}}$ through a dot-product operation followed by a sigmoid function. This process aims to capture the dynamic variations inherent in spatio-temporal representations. The outcome is then multiplied by the input representations and integrated via a residual connection, designed to activate the spatio-temporal information encapsulated within the feature embeddings. The output of the Temporal Activation module, denoted as $h_{ta}$, can be formulated as follows:
(8) $h_{ta}={\rm sigmoid}(\frac{{\rm FC}(e_{st})\cdot e_{l_{i}}^{T}}{\sqrt{d_{e_{l_{i}}}}})e_{l_{i}}+e_{l_{i}}.$
Employing the Temporal Target Attention mechanism to catch the interest preferences of different AOIs/Users towards Shops while simultaneously learning the variation across different times and spaces effectively. We achieve this by the linear projection of the Query, Key, and Value vectors with weights and bias (e.g., $[w_{Q},b_{Q}]$) generated by a fully connected neural network respectively. The module output denoted as $h_{tta}$, is articulated as follows:
(9) $h_{tta}={\rm softmax}(\frac{(w_{Q}\cdot e_{q}+b_{Q})(w_{K}\cdot e_{s}+b_{K})^{T}}{\sqrt{d_{k}}})(w_{V}\cdot e_{s}+b_{V}).$
Monotonic Response. To circumvent the difficulties in training when combining traditional linear monotonicity constraints with saturated (bounded) activation functions and the limitations that only convex models can be learned by combining convex activation functions, we construct a monotonic layer based on Constrained Monotonic Neural Networks (CMNN)(Runje and Shankaranarayana, 2023) framework. This approach not only facilitates utilizing non-saturated activation functions but also allows for a priori modeling of varying degrees of convexity and concavity across diverse application scenarios.
We employ a monotonicity indicator vector $t\in\{-1,0,1\}^{n}$ for the element-wise weights transformation of a feature vector to constrain the monotonicity. The operation $|\cdot|^{t}$ is applied as follows:
(10) $w_{j,i}^{\prime}=\begin{cases}\left|w_{j,i}\right|&\text{ if }t_{i}=1\\ -\left|w_{j,i}\right|&\text{ if }t_{i}=-1\\ w_{j,i}&\text{ otherwise }\end{cases}$
We rigorously define a convex activation function $\breve{\rho}\in\mathcal{\breve{A}}$, where $\mathcal{\breve{A}}$ denotes the set of all convex, monotonically increasing, lower-bounded, zero-centred functions.
Figure 3. CLU based activation functions construction
And we define the concave activation function $\hat{\rho}(x)$ and the saturated activation function $\tilde{\rho}(x)$ as follows:
(11) $\hat{\rho}(x)=-\breve{\rho}(-x)$
$\tilde{\rho}(x)=\begin{cases}\breve{\rho}(x+1)-\breve{\rho}(1)&\text{ if }x<0\\ \hat{\rho}(x-1)+\breve{\rho}(1)&\text{ otherwise }\end{cases}$
The function $\hat{\rho}(x)$ is concave and upper-bounded, while the function $\tilde{\rho}(x)$ is saturated and bounded. Furthermore, the activation selection vector $\mathbf{s}=(\breve{s},\hat{s},\tilde{s})\in\mathbb{N}^{3}$ and $\breve{s}+\hat{s}+\tilde{s}=m$ is introduced to approximate the concavity and convexity of the monotonic response through combining concave, convex, and saturated activation functions, where $m$ denotes the monotonic variables dimension.
Specifically, by setting $\mathbf{s}=(m,0,0)$, we guarantee that the monotonic response is modeled as a convex function, with analogous principles applied for concave modeling. Then the output of the combined activation functions $\rho^{\mathbf{s}}$ operates element-wise as follows:
(12) $\rho^{\mathbf{s}}(\mathbf{h})_{j}=\begin{cases}\breve{\rho}\left(h_{j}\right)&\text{ if }j\leq\breve{s}\\ \hat{\rho}\left(h_{j}\right)&\text{ if }\breve{s}<j\leq\breve{s}+\hat{s}\\ \tilde{\rho}\left(h_{j}\right)&\text{ otherwise }\end{cases}$
Simultaneously, to flexibly adapt to the spatiotemporally diverse property of marketing campaigns and learn the monotonic response functions, we introduce a strictly monotonic convex activation function, named Convex Linear Unit (CLU). Furthermore, we derive the corresponding concave activation function $\hat{\rho}(x)$ and the saturated activation function $\tilde{\rho}(x)$ through CLU. The activation functions construction is depicted in Figure 3, with CLU mathematical expressions as follows:
(13) $\breve{\rho}(x)=\begin{cases}-\frac{\omega_{0}}{2}+\frac{\omega_{0}}{1+e^{-\omega_{1}x}}&\text{ if }x<0\\ x&\text{ otherwise }\end{cases}$
The Universal Approximation Theorem(Cybenko, 1989; Hornik, 1991; Pinkus, 1999) posits that any continuous function over a closed interval can be approximated by neural networks with nonpolynomial activation functions. Concurrently, any multivariate continuous monotonic function on compact subsets of $\mathbb{R}^{k}$ can be approximated by monotonic constrained neural networks with at most $k$ hidden layers using sigmoid activation(Daniels and Velikova, 2010). Inspired by this, we propose the Four-Parameter Model (FPM) through an adaptive translation of the sigmoid function, combined with the CMNN framework, to flexibly approximate monotonic response functions in complex scenarios.
Figure 2(a) elucidates the functional form and parametric behavior of the FPM method, we learn and interpret the incentive sensitivity curve through four parameters, each bearing tangible significance in marketing. The parameter $\omega_{0}$ represents the natural conversion rate in the absence of any platform-driven incentive. The parameter $\omega_{1}$ denotes the sensitivity of the user at the motivational inflection point $\omega_{2}$, while $\omega_{3}$ corresponds to the maximal conversion rate attainable by the user.
Table 1. Overall performance on Exploding Red Packets and Delivery Fee Waiver.
Method Exploding Red Packets Delivery Fee Waiver
AUC $\uparrow$ MSE $\downarrow$ KL Div $\downarrow$ Corr Coeff $\uparrow$ MAE $\downarrow$ MSE $\downarrow$ KL Div $\downarrow$ Corr Coeff $\uparrow$
DNN(Schmidhuber, 2015) 0.6974 0.0636 0.6243 0.8513 0.1841 0.0831 0.8332 0.8304
DNN-M(Archer and Wang, 1993) 0.7022 0.0431 0.4259 0.8713 0.1232 0.0422 0.5347 0.8629
FPM 0.7113 0.0293 0.1901 0.9066 0.0913 0.0323 0.3083 0.8961
CMNN(Runje and Shankaranarayana, 2023)(ReLU(Nair and Hinton, 2010)) 0.7111 0.0282 0.1848 0.9150 0.0894 0.0298 0.2772 0.8916
CMNN(ELU(Clevert et al., 2015)) 0.7139 0.0273 0.1737 0.9171 0.0823 0.0281 0.2761 0.9127
CMNN(CLU) 0.7151 0.0269 0.1718 0.9189 0.0809 0.0274 0.2343 0.9143
CoMAN-B 0.7348 0.0224 0.1305 0.9229 0.0769 0.0263 0.2218 0.9193
CoMAN w/o AA 0.7414 0.0191 0.0984 0.9333 0.0624 0.0227 0.1897 0.9326
CoMAN w/o S-t 0.7407 0.0159 0.0925 0.9357 0.0733 0.0243 0.1809 0.9273
CoMAN 0.7480 0.0115 0.0901 0.9486 0.0454 0.0184 0.1211 0.9469
+0.0506 +0.0521 +0.5342 +0.0973 +0.1387 +0.0647 +0.7121 +0.1165
Table 2. Temporal segment evaluation on Exploding Red Packets. We denote the following periods: Bkfst (Breakfast), Lch (Lunch), A.Tea (Afternoon Tea), Din (Dinner), and M.S. (Midnight Snack).
Method MAE $\downarrow$ KL Div $\downarrow$
Bkfst Lch A.Tea Din M.S. Overall Bkfst Lch A.Tea Din M.S. Overall
FPM 0.1782 0.1385 0.1461 0.1297 0.1118 0.1325 0.1395 0.0824 0.1024 0.0918 0.0439 0.2801
CoMAN 0.0369 0.0449 0.0529 0.0780 0.0629 0.0448 0.0297 0.0288 0.0432 0.0737 0.0360 0.1480
+0.1413 +0.0936 +0.0932 +0.0517 +0.0489 +0.0877 +0.1098 +0.0536 +0.0592 +0.0181 +0.0079 +0.1321
Table 3. Temporal segment evaluation on Delivery Fee Waiver
Method MAE $\downarrow$ KL Div $\downarrow$
Bkfst Lch A.Tea Din M.S. Overall Bkfst Lch A.Tea Din M.S. Overall
FPM 0.1073 0.0883 0.1021 0.0932 0.0972 0.0943 0.0721 0.0809 0.0678 0.0733 0.0749 0.3692
CoMAN 0.0771 0.0569 0.0732 0.0631 0.0654 0.0632 0.0307 0.0276 0.0283 0.0259 0.0271 0.1398
+0.0302 +0.0314 +0.0289 +0.0301 +0.0318 +0.0311 +0.0414 +0.0533 +0.0395 +0.0474 +0.0478 +0.2294
Adaptive Response Enhancement. To enable the model to learn the variability of monotonic response across different spatio-temporal dimensions, we design the Adaptive Activation and Spatio-temporal Attention (S-t Attention) modules within the monotonic layer to enhance the model’s capability to perceive spatio-temporal information. These modules are designed to flexibly adapt the convexity, concavity, and parametric expressions of monotonic response functions through spatio-temporal representations.
In previous CMNN methodology, $\mathbf{s}$ is a priori parameter requiring manual adjustment through tuning rather than allowing for adaptive learning. The inflection point of a monotonic response function divides it into convex and concave segments. Drawing on these insights, we propose an Adaptive Activation module capable of adaptively learning and adjusting the proportion of convexity and concavity by catching and fusing deep representations with spatio-temporal information. As depicted on the left of the monotonic layer in Figure 1, we mask different parts of the monotonic representation $m_{i}$ with the adaptively learned convexity-concavity ratio factor $s_{st}$ and then combine through convex, concave, and saturated activation functions separately. The operations of the Adaptive Activation can be elucidated as follows:
(14) $s_{st}=|{\rm softmax}(\frac{m_{i}{({\rm Concat}(r_{f},e_{st})})^{T}}{\sqrt{d_{k}}})m_{i}|.$
(a) Amount
(b) Threshold
Figure 4. Incentive sensitivity monotonic response curve.
We design an Spatio-temporal (S-t) Attention module to utilize spatio-temporal information to further adaptively adjust and learn FPM. Firstly, we generate weights and bias ($[w_{i},b_{i}]$) terms for deep fused representation vectors (e.g., users, shops, etc.) through a fully connected layer. Then we conduct matrix multiplication calculations to derive the FPM parameters. The module structure is illustrated in Figure 2(b), and can be formulated as follows:
(15) $\omega_{i}={\rm FC}(w_{i}\cdot{\rm FC}(r_{f})+b_{i}).$
5. Experiments
5.1. Experimental Setup
(a) Overall
(b) Breakfast
(c) Lunch
(d) Afternoon Tea
(e) Dinner
(f) Midnight Snack
(g) City 1
(h) City 2
(i) City 3
Figure 5. Visualization of the prediction scores for incentive responses from various online models and the corresponding ground truth. Results are observed overall, with concurrent evaluations segmented by periods and cities.
Dataset Description. We conducted experiments on real-world datasets from two marketing campaigns on Ele.me, one of the largest OFOS platforms in China. The training datasets span one week, while the test dataset covers a single day. Notably, we randomly select 5% of online traffic flow as the detect traffic to randomly allocate incentive values as the ground truth for the incentive sensitivity function approximation. i) Exploding Red Packets Dataset: Send an average of 10 million exploding red packets to users online per day. It’s a classification dataset with approximately 1.05 million daily training samples and nearly 400 features, ultimately reaching 8.4 million samples. To evaluate the predictive accuracy of sensitivity functions, we trained models using unbiased samples collected from online detect traffic. ii) Delivery Fee Waiver Dataset: It is a regression dataset covering approximately 3.2 million AOIs, with about 4.3 million shops and 303 features. The daily samples are around 770 million, with the total exceeding 6 billion samples.
Evaluation Metric. To comprehensively evaluate the performance of our model, we assess it from both accuracy and distribution perspectives. We first employ metrics such as the Area Under the Curve (AUC)(Fawcett, 2006), Mean Absolute Error (MAE)(Willmott and Matsuura, 2005) and Mean Squared Error (MSE)(Hyndman and Koehler, 2006) to evaluate the precision of the model. Subsequently, we use the Kullback-Leibler Divergence (KL Div)(Kullback and Leibler, 1951) and Correlation Coefficients (Corr Coeff)(Pearson, 1895) to assess the similarity between the model’s predicted incentive sensitivity functions and the detect traffic (i.e. ground truth).
Baselines. To verify the effectiveness of our proposed methodology, we compare it against baselines with several state-of-the-art methods as follows: a) DNN: This is a naive model devoid of spatio-temporal and monotonic modeling capabilities. b) DNN-M: This method constrains the neural network weights to be exclusively non-negative or non-positive and combines sigmoid activation functions for monotonic modeling. c) FPM: This approach utilizes the mathematical four-parameter function to independently approximate the monotonic response function. d) CMNN: This method employs the CMNN framework in conjunction with different base convex activation functions, including CLU proposed in our study, while combining sigmoid as the final activation function.
Hyperparamters. In our experiments, for Exploding Red Packets, the model training batch size is set to 128 with Adagrad as an optimizer and a learning rate of 0.001. In Delivery Fee Waiver, the batch size is configured to 1024, using an Adagrad optimizer with a learning rate of 0.012. Furthermore, all experiments are conducted on AOP [^4] A model training platform self-developed by the Alibaba Group using 40 parameter servers and 400 worker threads.
5.2. Experimental Results
Ablation Study To comprehensively understand the capability of our proposed CoMAN model in approximating incentive sensitivity functions, and recognizing that the S-t attention and adaptive activation modules constitute the core modeling components of CoMAN, we carried out exhaustive ablation studies to evaluate the effectiveness of these pivotal modules. Accordingly, we introduced three ablation variants: a) CoMAN w/o (S-t and AA) (CoMAN-B), this baseline version employs CMNN based on CLU combined with FPM as the final activation function. b) CoMAN w/o AA, this revision excludes the Adaptive Activation module from the CoMAN model. c) CoMAN w/o S-t, this variant omits the S-t Attention module from the CoMAN structure. These versions are evaluated against the complete CoMAN model, which integrates all the modules, to systematically assess the impact of each module on the overall model performance. Specifically, the comparative experiments of CMNN without the spatio-temporal activation and the spatio-temporal target attention modules.
Meanwhile, we also utilize the CMNN framework to demonstrate the effectiveness of the proposed CLU activation function in adaptively adjusting incentive responses in diverse marketing campaigns. To validate the monotonic response property of the CoMAN, we conduct offline tests in which all non-monotonic features are held constant, while the monotonic features (e.g., amount and threshold) increase. Then construct the CVR monotonic response curves, as illustrated in Figures 4(a) and 4(b) respectively. To evaluate our monotonic modeling task, we present the output of the monotonic layer from offline experiments in Table 1, while Table 2 and Table 3 showcase the performance of the model during online marketing campaigns.
The results illustrated in the above tables indicate that the two modules within the CoMAN model’s monotonic layer align with our expectations, achieving significant improvements in both accuracy and prediction of real distributions. This demonstrates that CoMAN can adaptively adjust the convexity, concavity, and expression of the monotonic response functions in diverse marketing by enhancing the perception of spatio-temporal information. Moreover, as observed in Figure 5, we evaluate the model’s online predicted monotonic response curves both overall and across five periods, as well as for the top three cities with the most users. The fitting results for both the ablated and complete versions of CoMAN more closely approximate the ground truth distribution compared to the baseline, further verifying CoMAN’s effective capture of spatio-temporal characteristics in real-world marketing campaigns.
Online A/B Test We deploy our methodology to the Ele.me and conduct a two-week online A/B test. Compared with the online baseline model, our proposed model can bring more benefits while maintaining platform subsidies unchanged or even reduced, helping the marketing business improve budget efficiency and bring more growth. Detailed online outcomes as shown in Table 4.
Notably, within the regional pricing strategy for the delivery fee waiver marketing campaign, the CoMAN model also facilitates substantial business benefits for the platform. This is achieved even with a reduction in subsidy intensity by 0.76%, leading to a 0.70% increase in orders growth, a 0.62% improvement in the CTCVR, and a 1.06% increment in GMV.
Table 4. Online Business Performance Metrics. In comparison to the online FPM method, the CoMAN model demonstrated an improvement rate in the marketing of Ele.me.
Model CVR GMV Orders
CoMAN-B +0.82% +0.10% +0.45%
CoMAN w/o AA +1.20% +0.51% +1.56%
CoMAN w/o S-t +2.18% +3.02% +3.57%
CoMAN +3.09% +3.63% +3.68%
6. Conclusion
In this paper, we propose a novel constrained monotonic response model CoMAN for enhancing monotonic modeling by adaptive spatio-temporal awareness in diverse takeaway marketing. Specifically, we employ two spatio-temporal perception modules to activate and capture the spatio-temporal traits in the attribute representations. Subsequently, within the monotonic layer, we design two modules that utilize spatio-temporal information to enhance adaptive learning of concavity and convexity, as well as the sensitivity function expression. Comprehensive offline and online A/B testing conducted within two marketing campaigns on the Ele.me platform demonstrates the superiority of our proposed approach. Eventually, studies prove that our model more accurately reflects users’ incentive sensitivity across different locations and periods and exhibits superior pricing performance in various spatio-temporal dimensions, thus achieving the goal of improved marketing efficiency. Furthermore, future research will delve into the finer extraction of user sensitivity differences to incentives across disparate spatio-temporal dimensions.